OpenStudy (anonymous):

can sumone do the steps to this problem 5^−2∗3^−3∗4^0

OpenStudy (anonymous):

Is that: \[5^{-2}*3^{-3}*4^0\] If so: That's the same as: \[\frac{1}{5^{2}}*\frac{1}{3^{3}}*1\] \[\frac{1}{25}*\frac{1}{27}*1\] \[\frac{1}{675}\]